express the following as ratios having the same units in the numerator and denominator: 50 inches to 3 yards, 20 months to 3 years, 4 pounds to 42 ounces, 8 quarters to 12 dimes
Answer by stanbon(75887)

express the following as ratios having the same units in the numerator and denominator: 50 inches to 3 yards = 50in/3*36in = 25/3*18=25/54
-------------
20 months to 3 years=20 mts/3*12 mts
=5/3*3= 5/9
-----------------
4 pounds to 42 ounces= (4*16 ounces)/42 ounces
=32/21
------------
8 quarters to 12 dimes= 200 cents/120 cents = 5/3
-------------
Cheers,
Stan H.